# Evaluating machine learning models for predicting pesticides toxicity to honey bees

### v2

@@ Abstract
American Chemical Society,
@@ 1 Introduction
Low molecular weight compounds, often called small molecules, compounds, or just chemicals, play an important role in various research fields, including material science, biomedical research, and agrochemistry among others. Each field has distinct requirements for such molecules that can help define the chance of their success. These features are often related to physicochemical or molecular properties, for example, in biomedical research the Lipinski Rule of 5 is frequently used to assess lead and drug likeness 1, in agrochemistry similar physicochemical approaches were proposed for different types of pesticides 2, 3 and other types of crop protection compounds4. Similarly, important restraints have been established for compound toxicity, allowing both prediction 5, 6 and legislation considering harmful substances (e.g. REACH).
@@
The agrochemistry field shares similarities with biomedical research, often connected with evolutionary conserved mechanisms of actions or even molecular targets. However, significant differences between plants, insects, fungi, animals, and humans are frequently exploited when designing selective and specific small chemicals and biologics. Despite the abundance of medicinal chemistry datasets and benchmarks, such as MoleculeNet 7, PDBBind 8 and Therapeutics Data Commons (TDC) 9, curated agrochemical and environmental datasets remain scarce. This presents a significant challenge for developing robust predictive models. For example, while there are numerous models of human toxicity capable of predicting the off-target effects of known chemicals, there are few analogous models for agrochemicals 10. This gap is especially pronounced in the prediction of toxicity for environmentally important insects such as bees and specific water organisms, or economically important plants like crops.
@@
Large-scale worldwide use of herbicides and other pesticides can cause toxicities to non-target organisms and environmental degradation. A growing number of studies highlight the need for more selective herbicides targeting plant-specific pathways without affecting aquatic species, microbiota, insects, or mammals. 111213 Research on agrochemical toxicity has become even more essential with respect to environmentally significant insects like honey bee (Apis mellifera), which greatly influence both agriculture health and economies. This important pollinator, vulnerable to various chemicals, highlights both the necessity for more precise and species-specific toxicity data, and for more robust predictive models to assess and mitigate toxicity risks.
@@
The main scope of this research is the exploration of ApisTox, the most comprehensive dataset on bee toxicity 14 with various machine learning approaches to establish current predictive capabilities in this important domain. ApisTox includes chemical compounds represented as SMILES strings, along with their experimentally determined toxicity data for Apis mellifera. We assess the potential of ApisTox for the development of predictive methods ranging from molecular fingerprints to graph kernels, to graph neural networks, and pretrained neural models. Next, we analyze this dataset in terms of its uniqueness and similarities with respect to selected medicinal datasets included in the MoleculeNet benchmark. Finally, we draw conclusions on what works, what fails, and what types of deficiencies need to be addressed in order for the field of bee toxicity prediction to move forward. We explore the limits of current state-of-the-art machine learning (ML) methods, including comparisons with established methods such as molecular fingerprints or graph kernels. Explainable AI tools are also employed for behavioral testing, to verify the chemical soundness of predictions. s such, this work provides value not only to the ML community but also to experimental researchers and entities involved in toxicity studies, both from scientific and applied perspectives. Our study aims to contribute to ongoing efforts to reduce animal testing burdens and to support the development of safer agrochemicals, protect essential pollinators like bees, and promote sustainable agricultural practices.
@@ 2.1 ApisTox dataset
We focus on a recently published ApisTox dataset 14, which provides toxicity data on pesticides and other agrochemicals for the honey bee (Apis Mellifera). ApisTox consists of 1035 compounds, taken from the ECOTOX, PPDB, and BPDB databases, with an applied deduplication and standardization pipeline. This makes it larger and considerably higher quality than previous ML datasets in this area, such as BeeTOX 15. Notably, it contains a much higher number of 296 toxic molecules, as well as 739 non-toxic ones. The dataset is therefore moderately imbalanced, with 29% of pesticides being toxic. It is overall quite large, compared to most targeted agrochemical datasets and contains a very significant proportion of currently used pesticides. As such, it is highly representative of the overall pesticide chemical space while also providing annotations for the bee toxicity target.
Uniquely, ApisTox provides predetermined train-test splits, which ensures a fair comparison of different algorithms: maximum diversity split (MaxMin) 14, 16, as well as approximation of time split, based on PubChem literature reference. Those split approaches are shown to be potentially quite challenging for ML algorithms, with diverse test sets and high train-test separation, avoiding data leakage related to molecular similarity. This allows us to compare various ML-based molecular property prediction methods and extend this benchmark in the future.
@@ 2.2 Molecular dataset analysis methods
The original ApisTox publication 14 contains many analyzes, in particular showing that toxic and non-toxic pesticides vary in terms of molecular structure but not in basic physicochemical properties. In this work, we further analyze the properties of this dataset, showing that it covers a specific agrochemistry-related chemical space. This sets it apart from most benchmarks in molecular property prediction, which are focused on medicinal chemistry, e.g. MoleculeNet 7 and Therapeutics Data Commons (TDC) 9. In this section, we describe the chemoinformatics tools used for analysis of ApisTox and inter-dataset differences.
@@ 2.2.1 Molecular filters
Molecular filters are a set of rules used to select and limit a set of chemical molecules, which leads to obtaining the desired chemical space. They are used, among other things, in molecular screening to reduce the number of compounds to be analyzed by eliminating undesirable structures, such as those with low bioavailability or potential toxicity. Molecular filters can be divided into two main categories: structural filters, which classify molecules on the basis of the presence of specific functional groups or chemical fragments, and physicochemical filters, which evaluate compounds for selected properties such as molecular weight, logP, or the number of hydrogen bond donors and acceptors. A molecule passes the filter if it meets its criteria, e.g. does not contain any problematic groups, or has the properties in desired ranges. Molecular filter analysis allows the characterization and approximation of the types of chemical structures present in a given set of molecules 17.
Lipinski 1 and Ghose 18 filters were designed to identify drug-like compounds, focusing on parameters related to adsorption. Their criteria include molecular weight, number of hydrogen bonds, and solubility. Hao filter 2 aims to detect molecules with pesticide-specific properties, while the Tice Insecticides filter 3 is used to identify potential insecticides, taking into account the specific characteristics of this pesticide type. The Brenk filter 19 was developed for high-throughput screening (HTS). Specifically, it focuses on the elimination of compounds containing problematic functional groups that can exhibit high toxicity or adverse pharmacokinetic properties.
@@ 2.2.2 Molecular diversity
One of the simplest methods of analyzing such properties is comparing the distribution or types of chemical elements present in molecules, as well as substructures, functional groups, or fragments. A slight problem with such an analysis is that there is no single commonly accepted definition of a functional group or chemical fragment in the literature. Different approaches are usually used, based on different sets of functional groups, typically defined and published as collections of SMARTS patterns. Such methods are also often used as molecular fingerprints for building QSAR predictive models. Examples include the Laggner fingerprint 20 and MACCS Keys 21, but there are also more sophisticated approaches, like Ertl’s algorithm 22.
Arguably, the most popular method of comparing chemical spaces of datasets is checking the average pairwise Tanimoto of similarity between their molecules. Molecular fingerprints are typically used as a vectorization method, e.g. ECFP4 hashed fingerprint 23, or substructural PubChem fingerprint 24. The Tanimoto coefficient is considered the most suitable similarity measure for fingerprint-based molecular comparison 25.
ECFP4 fingerprint 23 is most commonly used as a vectorization method. It is an intuitive measure, with interpretable values. However, this approach has some downsides, such as being susceptible to clumping (clustering) of molecules 26, and is often only modestly correlated with biological activity 27.
#Circles measure 26, 27 has been proposed as a solution to these problems. Inspired by sphere exclusion clustering , it measures the coverage of the chemical space by identifying the local neighborhoods occupied by a given set of molecules. It counts the maximum number of mutually exclusive circles (hyperspheres) with a given radius in a dataset. Formally, given a set $\mathcal{S}$ with $n$ molecules, distance measure $d$ and threshold $t\in\left[0,1\right)$, we define #Circles measure as:
@@ 2.3.1 Molecular fingerprints
Descriptors, or descriptor sets, such as Mordred 28, compute a large, predefined set of molecular features, e.g. distribution of element types, topological indexes (e.g. Wiener, Zagreb), or substructure counts (e.g. rings, paths). They are highly interpretable and can achieve high performance when the target property is well correlated with physicochemical or topological features of the molecule.
Substructure fingerprints extract a predefined set of subgraphs, e.g. functional groups, aromatic rings, or atoms of a given element. The choice of substructures is typically selected by domain experts, most typically medicinal chemists. One can check the existence of a given substructure (binary fingerprint) or count its occurrences in a molecule (count fingerprint). Examples include MACCS Keys 21 and PubChem fingerprint 24. They are very interpretable and can encode domain-relevant features.
Hashed fingerprints define the general shape of subgraphs that are extracted from a molecule. Those can be, for example, circular neighborhoods for the ECFP fingerprint 23 or shortest paths between pairs of atoms in the Atom Pair fingerprint 29. A unique identifier is assigned to each subgraph, which is then hashed onto a constant-length output vector at a position based on the identifier value. In binary fingerprints, we ignore hashing collisions, whereas count fingerprints sum up all occurrences at a given index. They are the most flexible group, often resulting in very strong ML classifiers, competitive with complex GNNs 30, 31.
@@ 2.3.2 Feature engineering baselines
In one of the first papers on fair evaluation of GNNs for graph classification 32, the authors proposed to simply count atoms of different elements. This approach ignores the molecular graph topology, relying only on the distribution of the simplest atom attribute. However, it shows competitive performance on bioinformatics problems.
On the other hand, Local Topological Profile (LTP) 33 baseline uses graph topology exclusively. It creates a feature vector for a graph, based on concatenated histograms of topological descriptors: vertex degree, statistics of neighbors (bonded atoms) degrees (min, max, average, standard deviation), edge betweenness centrality (EBC), Jaccard Index (JI), and Local Degree Score (LDS). This approach aims to capture both local and global properties of graphs and is shown to be very competitive with complex GNNs.
Molecular Topological Profile (MOLTOP) 34 proposed a unified approach, with topological and molecular features, aimed specifically at molecular property prediction. First, histograms of topological descriptors are used, similarly to LTP, but different descriptors are selected, based on their chemoinformatical interpretation: EBC, Adjusted Rand Index (ARI), and SCAN Structural Similarity score. Then, simple statistics of basic molecular features are added to the feature vector: average, sum, and standard deviation of atom elements and bond orders. This fusion of two information sources is shown to give a particularly strong baseline for QSPR.
@@ 2.3.3 Graph kernels
Graph kernels quantify the similarity between graphs, based on their structure 35, 36, 37. In many cases, they also allow for incorporating vertex attributes, such as element type, and many molecule kernels have been proposed to describe the structural similarity of compounds. The resulting kernel matrix, containing pairwise similarity values between molecules, allows for the use of any established kernel methods, such as the SVM classifier. Due to their high expressive power, they show excellent performance in chemoinformatics 38, 39, 40.
The simplest possible kernels are the vertex and edge histogram kernels 41, which use a dot product between the counts of elements and the bonds between compounds, respectively. Many more complex kernels have been proposed, capturing local and global topology of graphs, e.g. shortest paths 42 or graphlets (all subgraphs up to a given size) 43, 44. The propagation kernel 45 has been proposed to diffuse the node information in the graph, strongly focusing on the functional similarity between molecules, in addition to their topology.
The Weisfeiler-Lehman (WL) kernel 46 is very commonly used, due to its excellent performance, speed, and strong theoretical properties. It is based on the Weisfeiler-Lehman isomorphism test 47, 48, aimed at distinguishing non-isomorphic graphs. The idea is to exchange information between neighboring nodes (bonded atoms), which captures the topological structure of a $k$-hop neighborhood around each atom after $k$ iterations. In particular, it has also directly influenced the Morgan algorithm for the molecular graph isomorphism test 49 and the ECFP fingerprint 23. This isomorphism testing framework also underlies the theory behind message-passing GNNs. The WL kernel is particularly computationally efficient, compared to most graph kernels.
The Weisfeiler-Lehman Optimal Assignment (WL-OA) kernel 50 has been proposed as an extension. The idea is that two similar molecules should have matching parts, e.g. same or similar rings or functional groups. This is realized by the optimal assignment kernel 51, which acts as a function mapping atoms in two molecules (identified using the Weisfeiler-Lehman procedure) to maximize the sum of the pairwise atom similarities. This substantially strengthens the ability of the kernel to distinguish between similar and dissimilar molecular substructures. Thus, it achieves very strong classification performance, albeit at a considerably higher computational cost.
@@ 2.3.4 Graph neural networks (GNNs)
Graph neural networks (GNNs) 52, 53, 54, 55 are a family of neural networks for graph representation learning, with appropriate inductive biases and theoretical guarantees. In particular, their operations, like graph convolution, are permutation-invariant, allowing them to process vertices and edges without assuming any natural order. They also naturally incorporate any atom and bond features. This makes them a natural choice for learning embeddings of molecules.
@@
Graph Convolutional Network (GCN) 56, based on regularized spectral graph convolution, takes a mean of the messages from the neighbors and the atom itself, weighting them by their degrees. GraphSAGE 57 has three variants, with “mean” variant being the most popular. It is quite similar to GCN, but first takes a simple mean of neighbors’ messages, and concatenates it to the vertex embedding before update. This separate treatment improves performance on heterophilous tasks 58, e.g. where the atom itself is much more important than its relationship to the neighborhood. The authors of Graph Isomorphism Network (GIN) 59 show that all message-passing GNNs are at most as powerful as the 1-WL test in distinguishing isomorphic graphs. Designed GIN convolution, based on multilayer perceptron (MLP) and sum pooling readout, is proven to be as powerful as WL test, resulting in good performance for graph classification. Graph Attention Network (GAT) 60 incorporates the attention mechanism to weight neighbors’ messages.
$k$-layer GNN is able to aggregate information about the $k$-hop neighborhood around each atom in its embedding. This is important for molecular classification since, e.g. understanding ring structures requires many layers under this paradigm. However, building deep GNNs has proven difficult, due to oversmoothing, oversquashing, and unstable training 61, which make it difficult to capture long-range dependencies. Jumping Knowledge 62 has been proposed as a simple solution, where we add skip connections, utilizing atom embeddings after each layer to obtain the molecule embedding, instead of just vectors after the final layer. Normalization layers, such as LayerNorm 63, 64, 65 can help stabilize training. Bayesian hyperparameter optimization (HPO) 66, e.g. using Parzen tree estimator 67, can be used to efficiently select good hyperparameter values, like learning rate or number of layers, enhancing the performance of the resulting architecture.
AttentiveFP 66 applied many of these ideas specifically for molecular property prediction. It utilizes the attention mechanism combined with artificial bonds for capturing long-range dependencies. It also uses the GRU recurrent network 68 for graph level readout, to focus on the most relevant atom embeddings. They also utilized Bayesian HPO to tune many hyperparameters of this architecture.

### v3

@@ Abstract
machine learning , chemoinformatics , molecular property prediction , graph neural networks , molecular fingerprints , ecotoxicology , agrochemistry
MSC:
92-04 , 92-08 , 92B20 92C99 , 92E10 , 68T10
journal: Computers in Biology and Medicine
@@ 1 Introduction
Low molecular weight compounds, often called small molecules, compounds, or just chemicals, play an important role in various research fields, including material science, biomedical research, and agrochemistry among others. Each field has distinct requirements for such molecules that can help define the chance of their success. These features are often related to physicochemical or molecular properties, for example, in biomedical research the Lipinski Rule of 5 is frequently used to assess lead and drug likeness [1], in agrochemistry similar physicochemical approaches were proposed for different types of pesticides [2, 3] and other types of crop protection compounds[4]. Similarly, important restraints have been established for compound toxicity, allowing both prediction [5, 6] and legislation considering harmful substances (e.g. REACH).
@@
The agrochemistry field shares similarities with biomedical research, often connected with evolutionary conserved mechanisms of actions or even molecular targets. However, significant differences between plants, insects, fungi, animals, and humans are frequently exploited when designing selective and specific small chemicals and biologics. Despite the abundance of medicinal chemistry datasets and benchmarks, such as MoleculeNet [7], PDBBind [8] and Therapeutics Data Commons (TDC) [9], curated agrochemical and environmental datasets remain scarce. This presents a significant challenge for developing robust predictive models. For example, while there are numerous models of human toxicity capable of predicting the off-target effects of known chemicals, there are few analogous models for agrochemicals [10]. This gap is especially pronounced in the prediction of toxicity for environmentally important insects such as bees and specific water organisms, or economically important plants like crops.
@@
Large-scale worldwide use of herbicides and other pesticides can cause toxicities to non-target organisms and environmental degradation. A growing number of studies highlight the need for more selective herbicides targeting plant-specific pathways without affecting aquatic species, microbiota, insects, or mammals. [11][12][13] Research on agrochemical toxicity has become even more essential with respect to environmentally significant insects like honey bee (Apis mellifera), which greatly influence both agriculture health and economies. This important pollinator, vulnerable to various chemicals, highlights both the necessity for more precise and species-specific toxicity data, and for more robust predictive models to assess and mitigate toxicity risks.
@@
The main scope of this research is the exploration of ApisTox, the most comprehensive dataset on bee toxicity [14] with various machine learning approaches to establish current predictive capabilities in this important domain. ApisTox includes chemical compounds represented as SMILES strings, along with their experimentally determined toxicity data for Apis mellifera. We assess the potential of ApisTox for the development of predictive methods ranging from molecular fingerprints to graph kernels, to graph neural networks, and pretrained neural models. Next, we analyze this dataset in terms of its uniqueness and similarities with respect to selected medicinal datasets included in the MoleculeNet benchmark. Finally, we draw conclusions on what works, what fails, and what types of deficiencies need to be addressed in order for the field of bee toxicity prediction to move forward. We explore the limits of current state-of-the-art machine learning (ML) methods, including comparisons with established methods such as molecular fingerprints or graph kernels. Explainable AI tools are also employed for behavioral testing, to verify the chemical soundness of predictions. s such, this work provides value not only to the ML community but also to experimental researchers and entities involved in toxicity studies, both from scientific and applied perspectives. Our study aims to contribute to ongoing efforts to reduce animal testing burdens and to support the development of safer agrochemicals, protect essential pollinators like bees, and promote sustainable agricultural practices.
@@ 2.1 ApisTox dataset
We focus on a recently published ApisTox dataset [14], which provides toxicity data on pesticides and other agrochemicals for the honey bee (Apis Mellifera). ApisTox consists of 1035 compounds, taken from the ECOTOX, PPDB, and BPDB databases, with an applied deduplication and standardization pipeline. This makes it larger and considerably higher quality than previous ML datasets in this area, such as BeeTOX [15]. Notably, it contains a much higher number of 296 toxic molecules, as well as 739 non-toxic ones. The dataset is therefore moderately imbalanced, with 29% of pesticides being toxic. It is overall quite large, compared to most targeted agrochemical datasets and contains a very significant proportion of currently used pesticides. As such, it is highly representative of the overall pesticide chemical space while also providing annotations for the bee toxicity target.
Uniquely, ApisTox provides predetermined train-test splits, which ensures a fair comparison of different algorithms: maximum diversity split (MaxMin) [14, 16], as well as approximation of time split, based on PubChem literature reference. Those split approaches are shown to be potentially quite challenging for ML algorithms, with diverse test sets and high train-test separation, avoiding data leakage related to molecular similarity. This allows us to compare various ML-based molecular property prediction methods and extend this benchmark in the future.
@@ 2.2 Molecular dataset analysis methods
The original ApisTox publication [14] contains many analyzes, in particular showing that toxic and non-toxic pesticides vary in terms of molecular structure but not in basic physicochemical properties. In this work, we further analyze the properties of this dataset, showing that it covers a specific agrochemistry-related chemical space. This sets it apart from most benchmarks in molecular property prediction, which are focused on medicinal chemistry, e.g. MoleculeNet [7] and Therapeutics Data Commons (TDC) [9]. In this section, we describe the chemoinformatics tools used for analysis of ApisTox and inter-dataset differences.
@@ 2.2.1 Molecular filters
Molecular filters are a set of rules used to select and limit a set of chemical molecules, which leads to obtaining the desired chemical space. They are used, among other things, in molecular screening to reduce the number of compounds to be analyzed by eliminating undesirable structures, such as those with low bioavailability or potential toxicity. Molecular filters can be divided into two main categories: structural filters, which classify molecules on the basis of the presence of specific functional groups or chemical fragments, and physicochemical filters, which evaluate compounds for selected properties such as molecular weight, logP, or the number of hydrogen bond donors and acceptors. A molecule passes the filter if it meets its criteria, e.g. does not contain any problematic groups, or has the properties in desired ranges. Molecular filter analysis allows the characterization and approximation of the types of chemical structures present in a given set of molecules [17].
Lipinski [1] and Ghose [18] filters were designed to identify drug-like compounds, focusing on parameters related to adsorption. Their criteria include molecular weight, number of hydrogen bonds, and solubility. Hao filter [2] aims to detect molecules with pesticide-specific properties, while the Tice Insecticides filter [3] is used to identify potential insecticides, taking into account the specific characteristics of this pesticide type. The Brenk filter [19] was developed for high-throughput screening (HTS). Specifically, it focuses on the elimination of compounds containing problematic functional groups that can exhibit high toxicity or adverse pharmacokinetic properties.
@@ 2.2.2 Molecular diversity
One of the simplest methods of analyzing such properties is comparing the distribution or types of chemical elements present in molecules, as well as substructures, functional groups, or fragments. A slight problem with such an analysis is that there is no single commonly accepted definition of a functional group or chemical fragment in the literature. Different approaches are usually used, based on different sets of functional groups, typically defined and published as collections of SMARTS patterns. Such methods are also often used as molecular fingerprints for building QSAR predictive models. Examples include the Laggner fingerprint [20] and MACCS Keys [21], but there are also more sophisticated approaches, like Ertl’s algorithm [22].
Arguably, the most popular method of comparing chemical spaces of datasets is checking the average pairwise Tanimoto of similarity between their molecules. Molecular fingerprints are typically used as a vectorization method, e.g. ECFP4 hashed fingerprint [23], or substructural PubChem fingerprint [24]. The Tanimoto coefficient is considered the most suitable similarity measure for fingerprint-based molecular comparison [25].
ECFP4 fingerprint [23] is most commonly used as a vectorization method. It is an intuitive measure, with interpretable values. However, this approach has some downsides, such as being susceptible to clumping (clustering) of molecules [26], and is often only modestly correlated with biological activity [27].
#Circles measure [26, 27] has been proposed as a solution to these problems. Inspired by sphere exclusion clustering , it measures the coverage of the chemical space by identifying the local neighborhoods occupied by a given set of molecules. It counts the maximum number of mutually exclusive circles (hyperspheres) with a given radius in a dataset. Formally, given a set $\mathcal{S}$ with $n$ molecules, distance measure $d$ and threshold $t\in\left[0,1\right)$, we define #Circles measure as:
@@ 2.3.1 Molecular fingerprints
Descriptors, or descriptor sets, such as Mordred [28], compute a large, predefined set of molecular features, e.g. distribution of element types, topological indexes (e.g. Wiener, Zagreb), or substructure counts (e.g. rings, paths). They are highly interpretable and can achieve high performance when the target property is well correlated with physicochemical or topological features of the molecule.
Substructure fingerprints extract a predefined set of subgraphs, e.g. functional groups, aromatic rings, or atoms of a given element. The choice of substructures is typically selected by domain experts, most typically medicinal chemists. One can check the existence of a given substructure (binary fingerprint) or count its occurrences in a molecule (count fingerprint). Examples include MACCS Keys [21] and PubChem fingerprint [24]. They are very interpretable and can encode domain-relevant features.
Hashed fingerprints define the general shape of subgraphs that are extracted from a molecule. Those can be, for example, circular neighborhoods for the ECFP fingerprint [23] or shortest paths between pairs of atoms in the Atom Pair fingerprint [29]. A unique identifier is assigned to each subgraph, which is then hashed onto a constant-length output vector at a position based on the identifier value. In binary fingerprints, we ignore hashing collisions, whereas count fingerprints sum up all occurrences at a given index. They are the most flexible group, often resulting in very strong ML classifiers, competitive with complex GNNs [30, 31].
@@ 2.3.2 Feature engineering baselines
In one of the first papers on fair evaluation of GNNs for graph classification [32], the authors proposed to simply count atoms of different elements. This approach ignores the molecular graph topology, relying only on the distribution of the simplest atom attribute. However, it shows competitive performance on bioinformatics problems.
On the other hand, Local Topological Profile (LTP) [33] baseline uses graph topology exclusively. It creates a feature vector for a graph, based on concatenated histograms of topological descriptors: vertex degree, statistics of neighbors (bonded atoms) degrees (min, max, average, standard deviation), edge betweenness centrality (EBC), Jaccard Index (JI), and Local Degree Score (LDS). This approach aims to capture both local and global properties of graphs and is shown to be very competitive with complex GNNs.
Molecular Topological Profile (MOLTOP) [34] proposed a unified approach, with topological and molecular features, aimed specifically at molecular property prediction. First, histograms of topological descriptors are used, similarly to LTP, but different descriptors are selected, based on their chemoinformatical interpretation: EBC, Adjusted Rand Index (ARI), and SCAN Structural Similarity score. Then, simple statistics of basic molecular features are added to the feature vector: average, sum, and standard deviation of atom elements and bond orders. This fusion of two information sources is shown to give a particularly strong baseline for QSPR.
@@ 2.3.3 Graph kernels
Graph kernels quantify the similarity between graphs, based on their structure [35, 36, 37]. In many cases, they also allow for incorporating vertex attributes, such as element type, and many molecule kernels have been proposed to describe the structural similarity of compounds. The resulting kernel matrix, containing pairwise similarity values between molecules, allows for the use of any established kernel methods, such as the SVM classifier. Due to their high expressive power, they show excellent performance in chemoinformatics [38, 39, 40].
The simplest possible kernels are the vertex and edge histogram kernels [41], which use a dot product between the counts of elements and the bonds between compounds, respectively. Many more complex kernels have been proposed, capturing local and global topology of graphs, e.g. shortest paths [42] or graphlets (all subgraphs up to a given size) [43, 44]. The propagation kernel [45] has been proposed to diffuse the node information in the graph, strongly focusing on the functional similarity between molecules, in addition to their topology.
The Weisfeiler-Lehman (WL) kernel [46] is very commonly used, due to its excellent performance, speed, and strong theoretical properties. It is based on the Weisfeiler-Lehman isomorphism test [47, 48], aimed at distinguishing non-isomorphic graphs. The idea is to exchange information between neighboring nodes (bonded atoms), which captures the topological structure of a $k$-hop neighborhood around each atom after $k$ iterations. In particular, it has also directly influenced the Morgan algorithm for the molecular graph isomorphism test [49] and the ECFP fingerprint [23]. This isomorphism testing framework also underlies the theory behind message-passing GNNs. The WL kernel is particularly computationally efficient, compared to most graph kernels.
The Weisfeiler-Lehman Optimal Assignment (WL-OA) kernel [50] has been proposed as an extension. The idea is that two similar molecules should have matching parts, e.g. same or similar rings or functional groups. This is realized by the optimal assignment kernel [51], which acts as a function mapping atoms in two molecules (identified using the Weisfeiler-Lehman procedure) to maximize the sum of the pairwise atom similarities. This substantially strengthens the ability of the kernel to distinguish between similar and dissimilar molecular substructures. Thus, it achieves very strong classification performance, albeit at a considerably higher computational cost.
@@ 2.3.4 Graph neural networks (GNNs)
Graph neural networks (GNNs) [52, 53, 54, 55] are a family of neural networks for graph representation learning, with appropriate inductive biases and theoretical guarantees. In particular, their operations, like graph convolution, are permutation-invariant, allowing them to process vertices and edges without assuming any natural order. They also naturally incorporate any atom and bond features. This makes them a natural choice for learning embeddings of molecules.
@@
Graph Convolutional Network (GCN) [56], based on regularized spectral graph convolution, takes a mean of the messages from the neighbors and the atom itself, weighting them by their degrees. GraphSAGE [57] has three variants, with “mean” variant being the most popular. It is quite similar to GCN, but first takes a simple mean of neighbors’ messages, and concatenates it to the vertex embedding before update. This separate treatment improves performance on heterophilous tasks [58], e.g. where the atom itself is much more important than its relationship to the neighborhood. The authors of Graph Isomorphism Network (GIN) [59] show that all message-passing GNNs are at most as powerful as the 1-WL test in distinguishing isomorphic graphs. Designed GIN convolution, based on multilayer perceptron (MLP) and sum pooling readout, is proven to be as powerful as WL test, resulting in good performance for graph classification. Graph Attention Network (GAT) [60] incorporates the attention mechanism to weight neighbors’ messages.
$k$-layer GNN is able to aggregate information about the $k$-hop neighborhood around each atom in its embedding. This is important for molecular classification since, e.g. understanding ring structures requires many layers under this paradigm. However, building deep GNNs has proven difficult, due to oversmoothing, oversquashing, and unstable training [61], which make it difficult to capture long-range dependencies. Jumping Knowledge [62] has been proposed as a simple solution, where we add skip connections, utilizing atom embeddings after each layer to obtain the molecule embedding, instead of just vectors after the final layer. Normalization layers, such as LayerNorm [63, 64, 65] can help stabilize training. Bayesian hyperparameter optimization (HPO) [66], e.g. using Parzen tree estimator [67], can be used to efficiently select good hyperparameter values, like learning rate or number of layers, enhancing the performance of the resulting architecture.
AttentiveFP [66] applied many of these ideas specifically for molecular property prediction. It utilizes the attention mechanism combined with artificial bonds for capturing long-range dependencies. It also uses the GRU recurrent network [68] for graph level readout, to focus on the most relevant atom embeddings. They also utilized Bayesian HPO to tune many hyperparameters of this architecture.